OpenStudy (anonymous):

Subtract. 12 11/14 and 2 9/14

OpenStudy (anonymous):

108/7

OpenStudy (anonymous):

its wrong.

OpenStudy (anonymous):

15 3/7

jhonyy9 (jhonyy9):

10 2/14 = 10 1/7

OpenStudy (anonymous):

Yes it's 10 1/7. I added both previously

OpenStudy (anonymous):

\[\left(12+\frac{11}{14}\right) -\left(2+\frac{9}{14}\right)=\frac{71}{7}=10\frac{1}{7} \]